# On generalized Beauville decompositions

### v1

@@ 2.6. Fourier transforms
    $(e^{\prime}_{\alpha},f^{\prime}_{\alpha}),e^{\prime}_{\beta}$ become $(e_{\sigma_{12}},f_{\sigma_{12}}),e_{\overline{\sigma}_{12}}$ respectively;
@@
    $e^{\prime}_{\overline{\Theta}},e^{\prime}_{H}$ become $e_{\sigma_{34}},e_{\overline{\sigma}_{34}}$ respectively.
Furthermore, we use (16) to define the operators $f^{\prime}_{\beta},f^{\prime}_{\overline{\Theta}},f^{\prime}_{H}$, completing the $\mathfrak{sl}_{2}$-triples with the raising operators $e^{\prime}_{\beta},e^{\prime}_{\overline{\Theta}},e^{\prime}_{H}$ respectively. For example
@@
Recall the operators (6). Combining Corollary 2.13, Lemma 2.10, and (15), we have
$e_{0}=[f^{\prime}_{\alpha},e^{\prime}_{\Theta}]=(-1)^{g}[f^{\prime}_{\alpha},e^{\prime}_{\overline{\Theta}}]+\mathrm{cst}\cdot[f^{\prime}_{\alpha},e^{\prime}_{\beta}].$
By Proposition 2.4(c) and (16), we obtain immediately that $[f^{\prime}_{\alpha},e^{\prime}_{\beta}]=0$. Therefore
(17) $e_{0}=(-1)^{g}[f^{\prime}_{\alpha},e^{\prime}_{\overline{\Theta}}].$
@@
$f_{\sigma_{12}}\mapsto f_{\sigma_{34}},\quad f_{\sigma_{34}}\mapsto-cf_{\sigma_{12}}+\mathrm{cst}\cdot f_{\overline{\sigma}_{34}}.$
@@
$\mathfrak{F}^{\mathfrak{g}}(f^{\prime}_{\alpha})=cf^{\prime}_{\overline{\Theta}},\quad\mathfrak{F}^{\mathfrak{g}}(f^{\prime}_{\overline{\Theta}})=-cf^{\prime}_{\alpha}+\mathrm{cst}\cdot f^{\prime}_{H}.$
@@
$f_{0}=-\mathfrak{F}^{\mathfrak{g}}(e_{0})$ $=(-1)^{g-1}[\mathfrak{F}^{\mathfrak{g}}(f^{\prime}_{\alpha}),\mathfrak{F}^{\mathfrak{g}}(e^{\prime}_{\overline{\Theta}})]$
$=(-1)^{g}[cf^{\prime}_{\overline{\Theta}},ce^{\prime}_{\alpha}]+\mathrm{cst}\cdot[f^{\prime}_{\overline{\Theta}},e^{\prime}_{H}]$
$=(-1)^{g}[f^{\prime}_{\overline{\Theta}},e^{\prime}_{\alpha}].$
Here the vanishing $[f^{\prime}_{\overline{\Theta}},e^{\prime}_{H}]=0$ follows again from Proposition 2.4(c) via (16). At this point, we have already shown that $f_{0}$ is of degree $-2$, since $f^{\prime}_{\overline{\Theta}},e^{\prime}_{\alpha}$ are of degrees 00, $-2$ respectively; more precisely
@@
$\mathfrak{F}^{\mathfrak{g}}(f_{0})$ $=(-1)^{g}[\mathfrak{F}^{\mathfrak{g}}(f^{\prime}_{\overline{\Theta}}),\mathfrak{F}^{\mathfrak{g}}(e^{\prime}_{\alpha})]$
$=(-1)^{g}[-cf^{\prime}_{\alpha},ce^{\prime}_{\overline{\Theta}}]+\mathrm{cst}\cdot[f^{\prime}_{\alpha},e^{\prime}_{\beta}]$
$=-(-1)^{g}[f^{\prime}_{\alpha},e^{\prime}_{\overline{\Theta}}]$
$=-e_{0}.$
@@
We have concluded that the decomposition induced by $(e_{0},h_{0},f_{0})$ is Fourier-stable. We normalize its indices to make the fundamental class $[\overline{J}_{C}]\in H^{0}(\overline{J}_{C},{\mathbb{C}})$ lie in $H_{(0)}^{0}(\overline{J}_{C},{\mathbb{C}})$. Since all three operators $e_{0},h_{0},f_{0}$ are rational, this gives rise to a decomposition with rational coefficients:
@@
Finally, we recall the following theorem from [41] concerning a canonical $\mathfrak{sl}_{2}$-triple associated with a general Lagrangian fibration $\pi:X\to B$.
@@
Let $\pi:X\to B$ be a Lagrangian fibration from a compact hyper-Kähler variety. Let $\Theta$ be a relatively ample class satisfying $q(\Theta)=0$ with respect to the BBF form $q(-)$. Then there is a unique $\mathfrak{sl}_{2}$-triple
@@
As before, we let $H$ be the pullback of a hyperplane class of the base $B$. By the proof of [41, Theorem 3.1], there is a Lie subalgebra
@@
whose raising operators are given by $e_{0}$ and $e_{1}:=H\cup-$ respectively, and the decomposition given by the first $\mathfrak{sl}_{2}$-triple $(e_{0},h_{0},f_{0})$ splits the perverse filtration. Furthermore, this decomposition is multiplicative with respect to the cup-product by [41, Proposition A.2]. This proves the existence part.
@@
In our setting $\pi:\overline{J}_{C}\to B$, since we have already proven that the operators (6) form an $\mathfrak{sl}_{2}$-triple with the desired degrees, this has to coincide with the $\mathfrak{sl}_{2}$-triple of Theorem 2.16. This completes the proofs of Theorems 0.3 and 0.5. ∎
@@ 3.1. Motivic Beauville decomposition and $\mathfrak{sl}_{2}$-triple
A natural framework for the motivic lifting is the theory of relative Chow motives of Corti–Hanamura [15]. It is built to be compatible with the decomposition theorem, adapts well to non-proper bases (e.g. the Hitchin fibration), and admits natural Chow/homological realizations [21]. We refer to [37, Chapter 8] and [33, Section 2.2] for a brief review of the theory.
@@
Here $\Delta_{\overline{J}_{C}/B}^{\mathrm{sm}}$ is the small relative diagonal of $\overline{J}_{C}$ in $\overline{J}_{C}\times_{B}\overline{J}_{C}\times_{B}\overline{J}_{C}$ and we refer to [33, Section 2.2.3] for the multiplicative structure of $\mathrm{CHM}(B)$.
@@
Note that the identities (19) together with $[\Delta_{\overline{J}_{C}/B}]=\sum_{i=0}^{2g}\mathfrak{p}_{i}$ already imply that the $\mathfrak{p}_{i}$ are orthogonal projectors. Indeed, we have
@@
(20) $[\Delta_{\overline{J}_{C}/B}^{\mathrm{sm}}]\circ\left((h_{0}+g\cdot[\Delta_{\overline{J}_{C}/B}])\times[\Delta_{\overline{J}_{C}/B}]+[\Delta_{\overline{J}_{C}/B}]\times(h_{0}+g\cdot[\Delta_{\overline{J}_{C}/B}])\right)\\ =(h_{0}+g\cdot[\Delta_{\overline{J}_{C}/B}])\circ[\Delta_{\overline{J}_{C}/B}^{\mathrm{sm}}].$
@@
When the base $B$ is proper the existence of $\Theta$ is well-studied in the context of compact hyper-Kähler geometry and is part of the Beauville–Voisin philosophy (see Section 3.3 below). Notably Rieß [40] proved that assuming the hyper-Kähler SYZ conjecture, any ${\mathbb{Q}}$-divisor $D$ with $q(D)=0$ on a compact hyper-Kähler variety $X$ of dimension $2n$ satisfies $D^{n+1}=0\in\mathrm{CH}^{n+1}(X,{\mathbb{Q}})$. Next, with the divisor $\Theta$ in (22) we define
@@ 3.2. Elliptic $K3$ surfaces
    $[\Delta^{\mathrm{sm}}_{S/\mathbb{P}^{1}}]\circ(h_{0}\times[\Delta_{S/\mathbb{P}^{1}}]+[\Delta_{S/\mathbb{P}^{1}}]\times h_{0}+[\Delta_{S/\mathbb{P}^{1}}]\times[\Delta_{S/\mathbb{P}^{1}}])=h_{0}\circ[\Delta^{\mathrm{sm}}_{S/\mathbb{P}^{1}}]$.
@@
In fact we have $\Delta_{S/{\mathbb{P}}^{1}*}\mathsf{s}=p_{1}^{*}\mathsf{s}\,p_{2}^{*}\mathsf{s}$ by definition. For $\Delta_{S/{\mathbb{P}}^{1}*}\mathsf{f}$ we may take $\mathsf{f}$ to be the class of a (singular) rational fiber of $\pi$. Since $[\Delta_{{\mathbb{P}}^{1}}]=[\mathrm{pt}\times{\mathbb{P}}^{1}]+[{\mathbb{P}}^{1}\times\mathrm{pt}]$, we find
@@
where the second to last equality uses $\Theta^{2}=0$. The other identity $[h_{0},f_{0}]=-2f_{0}$ and part (b) are both straightforward.
@@
where the second to last equality is because $-p_{1}^{*}\Theta\,p_{2}^{*}c+p_{1}^{*}c\,p_{2}^{*}\Theta$ is a 00-cycle of degree 00 supported on a chain of rational curves and hence vanishes. The third identity of (c) follows from the first two.
@@
$[\Delta_{S/{\mathbb{P}}^{1}}^{\mathrm{sm}}]-\left(q_{1}^{*}\Theta\,q_{23}^{*}[\Delta_{S/{\mathbb{P}}^{1}}]+\textrm{permutations}\right)+\left(q_{23}^{*}\Delta_{S/{\mathbb{P}}^{1}*}\Theta+\textrm{permutations}\right)=0.$
@@
(27) $[\Delta_{S/{\mathbb{P}}^{1}}^{\mathrm{sm}}]-\left(q_{1}^{*}\mathsf{s}\,q_{23}^{*}[\Delta_{S/{\mathbb{P}}^{1}}]+\textrm{permutations}\right)+(q_{2}^{*}\mathsf{s}\,q_{3}^{*}\mathsf{s}+\textrm{permutations})=0.$
@@
Next, we push the left-hand side of (27) all the way to the absolute triple product $S\times S\times S$, which sends $[\Delta_{S/{\mathbb{P}}^{1}}^{\mathrm{sm}}]$ to the small diagonal $[\Delta^{\mathrm{sm}}_{S}]\in\mathrm{CH}_{2}(S\times S\times S,{\mathbb{Q}})$. Using $\mathsf{s}\mathsf{f}=c$ and the identities
@@
$[\Delta_{S}^{\mathrm{sm}}]-(q_{1}^{\prime*}c\,q_{23}^{\prime*}[\Delta_{S}]+\textrm{permutations})+(q_{2}^{\prime*}c\,q_{3}^{\prime*}c+\textrm{permutations}),$
@@
As the left-hand side of (27) is a multiple of (28) whose pushforward to $S\times S\times S$ is clearly nonzero, this multiple must be 00. This proves the relation (27). ∎
@@ 3.3. Relations to Beauville–Voisin
Conjecture 3.6 is immediate for the elliptic $K3$ surface $\pi:S\to{\mathbb{P}}^{1}$ of Section 3.2 as $c\in\mathrm{CH}^{2}(h_{2}(S/{\mathbb{P}}^{1}),{\mathbb{Q}})$. The weaker, cohomological statement $c_{2i}(\overline{J}_{C})\in H^{4i}_{(2i)}(\overline{J}_{C},{\mathbb{Q}})$ holds for both the Beauville–Mukai system (and more generally for any Lagrangian fibration $\pi:X\to B$ from a compact hyper-Kähler variety by the calculations in [49, Introduction]) and the Hitchin system. In the Hitchin case, this can be seen as a consequence of $P=W$ and the natural splitting of the weight filtration on the Betti side.
@@
$e_{1}:=\Delta_{*}H\in\mathrm{Corr}^{1}(\overline{J}_{C},\overline{J}_{C})$
where $H\in\mathrm{CH}^{1}(\overline{J}_{C},{\mathbb{Q}})$ is the pullback of the hyperplane class of $B\simeq{\mathbb{P}}^{g}$, thus lifting $\mathfrak{sl}_{2}\times\mathfrak{sl}_{2}\subset\mathfrak{g}(\overline{J}_{C})$ (see [41]) to an action on the absolute Chow motive of $\overline{J}_{C}$. Further decompose the image of each $h_{i}(\overline{J}_{C}/B)$ into eigen-motives with respect to $h_{1}$. Finally, rearrange the eigen-motives with respect to $h:=h_{0}+h_{1}$, and use it to detect multiplicativity as well as the locations of the Chern classes.
@@
In this section we prove Theorem 0.4 using tautological relations on the universal Picard stack. For $g\geq 2$, let ¯$\overline{\mathcal{M}}_{g}$ be the moduli stack of stable curves of genus $g$. It is the Deligne–Mumford compactification of the moduli of genus $g$ nonsingular projective curves ${\mathcal{M}}_{g}\subset\overline{{\mathcal{M}}}_{g}$, and every curve on the boundary ¯$\overline{{\mathcal{M}}}_{g}\setminus{\mathcal{M}}_{g}$ has at worst simple nodes as singular points. More generally we also consider the Deligne–Mumford moduli stack of stable curves with marked points ${\mathcal{M}}_{g,n}\subset\overline{{\mathcal{M}}}_{g,n}$.
We denote by ¯$\overline{\mathcal{M}}_{g}^{\leq 1}\subset\overline{\mathcal{M}}_{g}$ the open locus where the curves are integral and have at most one node. Let ${\mathcal{C}}_{g}^{\leq 1}\to\overline{\mathcal{M}}_{g}^{\leq 1}$ be the universal curve with $\pi:\overline{J}_{g}^{\leq 1}\to\overline{\mathcal{M}}_{g}^{\leq 1}$ the corresponding compactified Jacobian fibration.
@@
    Let ¯$\overline{{\mathcal{M}}}^{\mathrm{int}}_{2}\subset\overline{{\mathcal{M}}}_{2}$ be the open locus of integral stable curves of genus $2$ with $\pi^{\mathrm{int}}:\overline{J}^{\mathrm{int}}_{2}\to\overline{{\mathcal{M}}}^{\mathrm{int}}_{2}$ the corresponding compactified Jacobian fibration. Then a generalized theta divisor does not exist for $\pi^{\mathrm{int}}$.

### v2

@@ 2.6. Fourier transforms
    $(e^{\prime}_{\alpha},f^{\prime}_{\alpha}),-e^{\prime}_{\beta}$ become $(e_{\sigma_{12}},f_{\sigma_{12}}),e_{\overline{\sigma}_{12}}$ respectively;
@@
    $e^{\prime}_{\overline{\Theta}},-c_{0}e^{\prime}_{\mathsf{H}}$ become $e_{\sigma_{34}},e_{\overline{\sigma}_{34}}$ respectively.
Furthermore, we use (17) to define the operators $f^{\prime}_{\beta},f^{\prime}_{\overline{\Theta}},f^{\prime}_{\mathsf{H}}$, completing the $\mathfrak{sl}_{2}$-triples with the raising operators $e^{\prime}_{\beta},e^{\prime}_{\overline{\Theta}},e^{\prime}_{\mathsf{H}}$ respectively. For example
@@
Recall the operators (6). Combining Corollary 2.14 with (14) and (16), we have
$e_{0}=[e_{\sigma},e^{\prime}_{\Theta}]=-[f^{\prime}_{\alpha},e^{\prime}_{\Theta}]=-[f^{\prime}_{\alpha},-c_{0}e^{\prime}_{\overline{\Theta}}]+\mathrm{cst}\cdot[f^{\prime}_{\alpha},e^{\prime}_{\beta}].$
By Proposition 2.4(c) and (17), we obtain immediately that $f^{\prime}_{\alpha},e^{\prime}_{\beta}]=0$. Therefore
(18) $e_{0}=c_{0}[f^{\prime}_{\alpha},e^{\prime}_{\overline{\Theta}}].$
@@
$f_{\sigma_{12}}\mapsto f_{\sigma_{34}},\quad f_{\sigma_{34}}\mapsto-f_{\sigma_{12}}+\mathrm{cst}\cdot f_{\overline{\sigma}_{34}}.$
@@
$\mathfrak{F}^{\mathfrak{g}}(f^{\prime}_{\alpha})=c_{1}f^{\prime}_{\overline{\Theta}},\quad\mathfrak{F}^{\mathfrak{g}}(f^{\prime}_{\overline{\Theta}})=-c_{1}f^{\prime}_{\alpha}+\mathrm{cst}\cdot f^{\prime}_{\mathsf{H}}.$
@@
$f_{0}=-\mathfrak{F}^{\mathfrak{g}}(e_{0})$ $=-c_{0}[\mathfrak{F}^{\mathfrak{g}}(f^{\prime}_{\alpha}),\mathfrak{F}^{\mathfrak{g}}(e^{\prime}_{\overline{\Theta}})]$
$=-c_{0}[c_{1}f^{\prime}_{\overline{\Theta}},-c_{1}e^{\prime}_{\alpha}]+\mathrm{cst}\cdot[f^{\prime}_{\overline{\Theta}},e^{\prime}_{\mathsf{H}}]$
$=c_{0}[f^{\prime}_{\overline{\Theta}},e^{\prime}_{\alpha}].$
Here the vanishing $f^{\prime}_{\overline{\Theta}},e^{\prime}_{\mathsf{H}}]=0$ follows again from Proposition 2.4(c) via (17). At this point, we have already shown that $f_{0}$ is of degree 2-2, since $f^{\prime}_{\overline{\Theta}},e^{\prime}_{\alpha}$ are of degrees 0, 2-2 respectively; more precisely
@@
$\mathfrak{F}^{\mathfrak{g}}(f_{0})$ $=c_{0}[\mathfrak{F}^{\mathfrak{g}}(f^{\prime}_{\overline{\Theta}}),\mathfrak{F}^{\mathfrak{g}}(e^{\prime}_{\alpha})]$
$=c_{0}[-c_{1}f^{\prime}_{\alpha},c_{1}e^{\prime}_{\overline{\Theta}}]+\mathrm{cst}\cdot[f^{\prime}_{\alpha},e^{\prime}_{\beta}]$
$=-c_{0}[f^{\prime}_{\alpha},e^{\prime}_{\overline{\Theta}}]=-e_{0}.$
@@
We have concluded that the decomposition induced by $(e_{0},h_{0},f_{0})$ is Fourier-stable. We normalize its indices to make the fundamental class $\overline{J}_{C}]\in H^{0}(\overline{J}_{C},{\mathbb{C}})$ lie in $H_{(0)}^{0}(\overline{J}_{C},{\mathbb{C}})$. Since all three operators $e_{0},h_{0},f_{0}$ are rational, this gives rise to a decomposition with rational coefficients:
@@
Finally, we recall the following theorem concerning a canonical $\mathfrak{sl}_{2}$-triple associated with a general Lagrangian fibration $\pi:X\to B$. It summarizes several results from [43].
@@
Let $\pi:X\to B$ be a Lagrangian fibration from a compact hyper-Kähler variety. Let $\Theta$ be a relatively ample class satisfying $q(\Theta)=0$ with respect to the BBF form q()q(-). Then there is a unique $\mathfrak{sl}_{2}$-triple
@@
As before, we let $\mathsf{H}$ be the pullback of a hyperplane class of the base $B$. By the proof of [43, Theorem 3.1], there is a Lie subalgebra
@@
whose raising operators are given by $e_{0}$ and $e_{1}:=\mathsf{H}\cup$- respectively, and the decomposition given by the first $\mathfrak{sl}_{2}$-triple $(e_{0},h_{0},f_{0})$ splits the perverse filtration. Furthermore, this decomposition is multiplicative with respect to the cup-product by [43, Proposition A.2]. This proves the existence part.
@@
In our setting $\pi:\overline{J}_{C}\to B$, since we have already proven that the operators (6) form an $\mathfrak{sl}_{2}$-triple with the desired degrees, this has to coincide with the canonical $\mathfrak{sl}_{2}$-triple of Theorem 2.17. This completes the proofs of Theorems 0.3 and 0.5. ∎
@@ 3.1. Motivic Beauville decomposition and $\mathfrak{sl}_{2}$-triple
A natural framework for the motivic lifting is the theory of relative Chow motives of Corti–Hanamura [16]. It is built to be compatible with the decomposition theorem, adapts well to non-proper bases (e.g. the Hitchin fibration), and admits natural Chow/homological realizations [22]. We refer to [39, Chapter 8] and [34, Section 2.2] for a brief review of the theory.
@@
Here $\Delta_{\overline{J}_{C}/B}^{\mathrm{sm}}$ is the small relative diagonal of $\overline{J}_{C}$ in $\overline{J}_{C}\times_{B}\overline{J}_{C}\times_{B}\overline{J}_{C}$ and we refer to [34, Section 2.2.3] for the multiplicative structure of $\mathrm{CHM}(B)$.
@@
Note that the identities (20) together with $\Delta_{\overline{J}_{C}/B}]=\sum_{i=0}^{2g}\mathfrak{p}_{i}$ already imply that the $\mathfrak{p}_{i}$ are orthogonal projectors. Indeed, we have
@@
(21) $\Delta_{\overline{J}_{C}/B}^{\mathrm{sm}}]\circ\left((h_{0}+g\cdot[\Delta_{\overline{J}_{C}/B}])\times[\Delta_{\overline{J}_{C}/B}]+[\Delta_{\overline{J}_{C}/B}]\times(h_{0}+g\cdot[\Delta_{\overline{J}_{C}/B}])\right)\\ =(h_{0}+g\cdot[\Delta_{\overline{J}_{C}/B}])\circ[\Delta_{\overline{J}_{C}/B}^{\mathrm{sm}}].$
@@
When the base $B$ is proper the existence of $\Theta$ is well-studied in the context of compact hyper-Kähler geometry and is part of the Beauville–Voisin philosophy (see Section 3.3 below). Notably Rieß [42] proved that assuming the hyper-Kähler SYZ conjecture, any $\mathbb{Q}$}-divisor $D$ with $q(D)=0$ on a compact hyper-Kähler variety $X$ of dimension $2n$ satisfies
$D^{n+1}=0\in\mathrm{CH}^{n+1}(X,{\mathbb{Q}}).$
Next, with the divisor $\Theta$ in (23) we define
@@ 3.2. Elliptic $K3$ surfaces
    $\Delta^{\mathrm{sm}}_{S/\mathbb{P}^{1}}]\circ(h_{0}\times[\Delta_{S/\mathbb{P}^{1}}]+[\Delta_{S/\mathbb{P}^{1}}]\times h_{0}+[\Delta_{S/\mathbb{P}^{1}}]\times[\Delta_{S/\mathbb{P}^{1}}])=h_{0}\circ[\Delta^{\mathrm{sm}}_{S/\mathbb{P}^{1}}]$.
@@
In fact we have $\Delta_{S/{\mathbb{P}}^{1}*}\mathsf{s}=p_{1}^{*}\mathsf{s}\,p_{2}^{*}\mathsf{s}$ by definition. For $\Delta_{S/{\mathbb{P}}^{1}*}\mathsf{f}$ we may take $\mathsf{f}$ to be the class of a (singular) rational fiber of $\pi$. Since $\Delta_{{\mathbb{P}}^{1}}]=[\mathrm{pt}\times{\mathbb{P}}^{1}]+[{\mathbb{P}}^{1}\times\mathrm{pt}]$, we find
@@
where the second to last equality uses $\Theta^{2}=0$. The other identity $h_{0},f_{0}]=-2f_{0}$ and part (b) are both straightforward.
@@
where the second to last equality is because $p_{1}^{*}\Theta\,p_{2}^{*}c+p_{1}^{*}c\,p_{2}^{*}\Theta$ is a 0-cycle of degree 0 supported on a chain of rational curves and hence vanishes. The third identity of (c) follows from the first two.
@@
$\Delta_{S/{\mathbb{P}}^{1}}^{\mathrm{sm}}]-\left(q_{1}^{*}\Theta\,q_{23}^{*}[\Delta_{S/{\mathbb{P}}^{1}}]+\textrm{permutations}\right)+\left(q_{23}^{*}\Delta_{S/{\mathbb{P}}^{1}*}\Theta+\textrm{permutations}\right)=0.$
@@
(28) $\Delta_{S/{\mathbb{P}}^{1}}^{\mathrm{sm}}]-\left(q_{1}^{*}\mathsf{s}\,q_{23}^{*}[\Delta_{S/{\mathbb{P}}^{1}}]+\textrm{permutations}\right)+(q_{2}^{*}\mathsf{s}\,q_{3}^{*}\mathsf{s}+\textrm{permutations})=0.$
@@
Next, we push the left-hand side of (28) all the way to the absolute triple product $S\times S\times S$, which sends $\Delta_{S/{\mathbb{P}}^{1}}^{\mathrm{sm}}]$ to the small diagonal $\Delta^{\mathrm{sm}}_{S}]\in\mathrm{CH}_{2}(S\times S\times S,{\mathbb{Q}})$. Using $\mathsf{s}\mathsf{f}=c$ and the identities
@@
$\Delta_{S}^{\mathrm{sm}}]-(q_{1}^{\prime*}c\,q_{23}^{\prime*}[\Delta_{S}]+\textrm{permutations})+(q_{2}^{\prime*}c\,q_{3}^{\prime*}c+\textrm{permutations}),$
@@
As the left-hand side of (28) is a multiple of (29) whose pushforward to $S\times S\times S$ is clearly nonzero, this multiple must be 0. This proves the relation (28). ∎
@@ 3.3. Relations to Beauville–Voisin
Conjecture 3.6 is immediate for the elliptic $K3$ surface $\pi:S\to{\mathbb{P}}^{1}$ of Section 3.2 as $c\in\mathrm{CH}^{2}(h_{2}(S/{\mathbb{P}}^{1}),{\mathbb{Q}})$. The weaker, cohomological statement $c_{2i}(\overline{J}_{C})\in H^{4i}_{(2i)}(\overline{J}_{C},{\mathbb{Q}})$ holds for both the Beauville–Mukai system (and more generally for any Lagrangian fibration $\pi:X\to B$ from a compact hyper-Kähler variety by the calculations in [51, Introduction]) and the Hitchin system. In the Hitchin case, this can be seen as a consequence of $P=W$ and the natural splitting of the weight filtration on the Betti side.
@@
$e_{1}:=\Delta_{*}\mathsf{H}\in\mathrm{Corr}^{1}(\overline{J}_{C},\overline{J}_{C})$
where $\mathsf{H}\in\mathrm{CH}^{1}(\overline{J}_{C},{\mathbb{Q}})$ is the pullback of the hyperplane class of $B\simeq{\mathbb{P}}^{g}$, thus lifting $\mathfrak{sl}_{2}\times\mathfrak{sl}_{2}\subset\mathfrak{g}(\overline{J}_{C})$ (see [43]) to an action on the absolute Chow motive of $\overline{J}_{C}$. Further decompose the image of each $h_{i}(\overline{J}_{C}/B)$ into eigen-motives with respect to $h_{1}$. Finally, rearrange the eigen-motives with respect to $h:=h_{0}+h_{1}$, and use it to detect multiplicativity as well as the locations of the Chern classes.
@@
In this section we prove Theorem 0.4 using tautological relations on the universal Picard stack. For $g\geq 2$, let $\overline{\mathcal{M}}_{g}$ be the moduli stack of stable curves of genus $g$. It is the Deligne–Mumford compactification of the moduli stack of genus $g$ nonsingular projective curves ${\mathcal{M}}_{g}\subset\overline{{\mathcal{M}}}_{g}$, and every curve on the boundary $\overline{{\mathcal{M}}}_{g}\setminus{\mathcal{M}}_{g}$ has at worst simple nodes as singular points. More generally we also consider the Deligne–Mumford moduli stack of stable curves with marked points ${\mathcal{M}}_{g,n}\subset\overline{{\mathcal{M}}}_{g,n}$.
We denote by $\overline{\mathcal{M}}_{g}^{\leq 1}\subset\overline{\mathcal{M}}_{g}$ the open locus where the curves are integral and have at most one node. Let ${\mathcal{C}}_{g}^{\leq 1}\to\overline{\mathcal{M}}_{g}^{\leq 1}$ be the universal curve with $\pi:\overline{J}_{g}^{\leq 1}\to\overline{\mathcal{M}}_{g}^{\leq 1}$ the corresponding compactified Jacobian fibration.
@@
    Let $\overline{{\mathcal{M}}}^{\mathrm{int}}_{2}\subset\overline{{\mathcal{M}}}_{2}$ be the open locus of integral stable curves of genus $2$ with $\pi^{\mathrm{int}}:\overline{J}^{\mathrm{int}}_{2}\to\overline{{\mathcal{M}}}^{\mathrm{int}}_{2}$ the corresponding compactified Jacobian fibration. Then a generalized theta divisor does not exist for $\pi^{\mathrm{int}}$.